OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5

hero (hero):

-3(x + 3) = -3(x + 1) - 5 -3(x + 1 + 2) = -3(x + 1) - 5 -3(x + 1) - 6 = -3(x + 1) - 5 3(x + 1) - 3(x + 1) = 6 - 5 0 = 1 False

hero (hero):

No solution

OpenStudy (anonymous):

thanks